### v2

@@ 5.2. General Lie algebras
We will now show that every nilpotent Lie algebra is mm–great for sufficiently large $m$. Our approach will largely mirror that of the proofs of Propositions 5.1 and 5.2, but taking advantage of the larger value of $m$.
@@
Combining Proposition 5.5 with Theorem 3.7 gives us Theorem A. We also show that $m$-greatness is preserved by taking products and factors.
@@ 5.3. Counter Example
Although $m$-greatness is a useful property, it has some weaknesses. In particular, although it is expected that on any nilmanifold almost any random walk generated by two translations will be rapid mixing, not all Lie algebras are 22–great. We provide an explicit example of such a Lie algebra.
@@
(5.5) $\begin{split}[X_{1},X_{2}]=Y_{1},\quad[X_{1},X_{3}]=Y_{2},\quad[X_{2},X_{3}]=Y_{3},\\ [Y_{1},X_{1}]=Z_{1},\quad[Y_{1},X_{2}]=Z_{2},\quad[Y_{1},X_{3}]=Z_{3},\\ [Y_{2},X_{1}]=Z_{4},\quad[Y_{2},X_{2}]=Z_{5},\quad[Y_{2},X_{3}]=Z_{6},\\ [Y_{3},X_{1}]=Z_{5}-Z_{3},\quad[Y_{3},X_{2}]=Z_{7},\quad[Y_{3},X_{3}]=Z_{8},\\ [Z_{1},X_{3}]=3W,\quad[Z_{2},X_{3}]=-3W,\quad[Z_{3},X_{1}]=-W,\\ [Z_{3},X_{2}]=W,\quad[Z_{3},X_{3}]=-2W,\quad[Z_{4},X_{2}]=-3W,\\ [Z_{5},X_{1}]=W,\quad[Z_{5},X_{2}]=2W,\quad[Z_{5},X_{3}]=-W,\\ [Z_{6},X_{2}]=3W,\quad[Z_{7},X_{1}]=-3W,\quad[Z_{8},X_{1}]=-3W,\\ [Y_{1},Y_{2}]=4W,\quad[Y_{1},Y_{3}]=-4W,\quad[Y_{2},Y_{3}]=-4W,\end{split}$
@@ Appendix A Lemma for application of Diophantine estimate
Note that the singularity of the RHS is removable since it can be rewritten as $g(\theta):=2\sum_{n=0}^{\infty}\frac{(\pi\theta)^{2n}}{(2n+2)!}$. Since $g$ is positive and continuous on [1/2,1/2][-1/2,1/2] its achieves its minimum value $C.$
@@ Appendix B Central Limit Theorem
Let $B=(1-{\mathcal{L}})^{-1}A=\sum_{n=0}^{\infty}{\mathcal{L}}^{n}A.$ By (3.2) this series converges in $C^{0}$ for $r$ large enough. Thus $B(x_{n})=A(x_{n})+{\mathbb{E}}(B(x_{n+1}|{\mathcal{F}}_{n})$ where ${\mathcal{F}}_{n}$ is the $\sigma$-algebra generated by $(x_{0},\dots,x_{n}).$ Summing up, we get $\sum_{n=0}^{N-1}A(x_{n})\!\!=\!\!\sum_{n=1}^{N}\Delta_{n}\!\!+\!\!B(x_{0})\!\!-\!\!B(x_{N})$ where $\Delta_{n}\!\!=\!\!B(x_{n})\!-\!{\mathbb{E}}(B(x_{n})|{\mathcal{F}}_{n-1})$ is a martingale difference sequence. Now by the CLT for martingales (see e.g. [8, §8.2]), to prove the CLT, it is sufficient to show that the limit $\sigma^{2}:=\lim_{N\to\infty}\frac{1}{N}\sum_{n=0}^{N-1}q_{n}$ exists (in probability) where $q_{n}={\mathbb{E}}(\Delta_{n}^{2}|{\mathcal{F}}_{n-1}).$ Note that in our case $q_{n}=Q(x_{n-1})$ for a continuous function $Q$ so the existence of the limit follows from the ergodicity of our Markov chain. ∎

### v3

@@ 5.2. General Lie algebras
We will now show that every nilpotent Lie algebra is mm–great for sufficiently large $m$. Our approach will largely mirror that of the proofs of but taking advantage of the larger value of $m$.
@@
Combining Proposition 5.5 with Theorem 3.7 gives us Theorem A.
We also note that $m$-greatness is preserved by taking products and factors.
@@ 5.3. Counter Example
Although $m$-greatness is a great property, it has some weaknesses. In particular, although it is expected that on any nilmanifold almost any random walk generated by two translations will be rapid mixing, not all Lie algebras are 22–great. We provide an explicit example of such a Lie algebra.
@@
$\begin{split}[X_{1},X_{2}]=Y_{1},\quad[X_{1},X_{3}]=Y_{2},\quad[X_{2},X_{3}]=Y_{3},\\ [Y_{1},X_{1}]=Z_{1},\quad[Y_{1},X_{2}]=Z_{2},\quad[Y_{1},X_{3}]=Z_{3},\\ [Y_{2},X_{1}]=Z_{4},\quad[Y_{2},X_{2}]=Z_{5},\quad[Y_{2},X_{3}]=Z_{6},\\ [Y_{3},X_{1}]=Z_{5}-Z_{3},\quad[Y_{3},X_{2}]=Z_{7},\quad[Y_{3},X_{3}]=Z_{8},\\ [Z_{1},X_{3}]=3W,\quad[Z_{2},X_{3}]=-3W,\quad[Z_{3},X_{1}]=-W,\\ [Z_{3},X_{2}]=W,\quad[Z_{3},X_{3}]=-2W,\quad[Z_{4},X_{2}]=-3W,\\ [Z_{5},X_{1}]=W,\quad[Z_{5},X_{2}]=2W,\quad[Z_{5},X_{3}]=-W,\\ [Z_{6},X_{2}]=3W,\quad[Z_{7},X_{1}]=-3W,\quad[Z_{8},X_{1}]=-3W,\\ [Y_{1},Y_{2}]=4W,\quad[Y_{1},Y_{3}]=-4W,\quad[Y_{2},Y_{3}]=-4W,\end{split}$
@@ Appendix A Lemma for application of Diophantine estimate
The singularity of the RHS is removable since it can be rewritten as $g(\theta)\!\!:=\!2\!\!\sum_{n=0}^{\infty}\frac{(\pi\theta)^{2n}}{(2n+2)!}$. Since $g$ is positive and continuous on [1/2,1/2][-1/2,1/2] its achieves its minimum value $C.$
@@ Appendix B Central Limit Theorem
Let $B=(1-{\mathcal{L}})^{-1}A=\sum_{n=0}^{\infty}{\mathcal{L}}^{n}A.$ By (3.2) this series converges in $C^{0}$ for $r$ large enough. Thus $B(x_{n})=A(x_{n})+{\mathbb{E}}(B(x_{n+1}|{\mathcal{F}}_{n})$ where ${\mathcal{F}}_{n}$ is the $\sigma$-algebra generated by $(x_{0},\dots,x_{n}).$ Summing up, we obtain $\sum_{n=0}^{N-1}A(x_{n})=\sum_{n=1}^{N}\Delta_{n}+B(x_{0})-B(x_{N})$ where $\Delta_{n}\!\!=\!\!B(x_{n})\!-\!{\mathbb{E}}(B(x_{n})|{\mathcal{F}}_{n-1})$ is a martingale difference sequence. Now by the CLT for martingales (see e.g. [8, §8.2]), to prove the CLT, it is sufficient to show that the limit $\sigma^{2}:=\lim_{N\to\infty}\frac{1}{N}\sum_{n=0}^{N-1}q_{n}$ exists (in probability) where $q_{n}={\mathbb{E}}(\Delta_{n}^{2}|{\mathcal{F}}_{n-1}).$ Note that in our case $q_{n}=Q(x_{n-1})$ for a continuous function $Q$ so the existence of the limit follows from the ergodicity of our Markov chain. ∎
@@ Appendix C Lie algebra calculation
Statements and Declarations
The authors have no relevant financial or non-financial interests to disclose. Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.